OpenStudy (anonymous):

√(6x+5)= √53 Show Work

OpenStudy (anonymous):

(6x+5)^(1/2) = (53)^(1/2) square both the side 6x+5 = 53 6x = 48 x = 48/6 x = 8

OpenStudy (anonymous):

√(6x+5)= √53 square both sides to get 6x + 5 = 53 6x = 53 - 5 x = 48/6 x = 8

OpenStudy (anonymous):

square both sides 6x+5=53 subtract 5 from both sides 6x=48 divide both sides by 6 x=8